# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ 2.2. A quasi-convexity characterization of $C^{1,1}$
$\eta_{\varepsilon}\vcentcolon=\frac{1}{\varepsilon^{n}}\,\eta\!\left(\frac{\cdot}{\varepsilon}\right),\quad\forall\varepsilon>0.$
@@
Suppose $u$ is $\lambda$-quasi-convex on ${\mathbb{R}}^{n}$ and let $\eta$ be a mollifier; let $\{\eta_{\varepsilon}\}_{\varepsilon>0}$ the approximate identity based on $\eta$. Then $u_{\varepsilon}\vcentcolon=u\ast\eta_{\varepsilon}$ is $\lambda$-quasi-convex.
@@
Let $f\vcentcolon=u+\frac{\lambda}{2}|\cdot|^{2}$ be convex. It is easy to see that $f_{\varepsilon}\vcentcolon=f\ast\eta_{\varepsilon}$ is still convex and is smooth. In addition one has
$\begin{split}\left(|\cdot|^{2}\ast\eta_{\varepsilon}\right)\!(x)\vcentcolon=\int_{{\mathbb{R}}^{n}}|x-z|^{2}\eta(z/\varepsilon)\varepsilon^{-n}\mathrm{d}z=|x|^{2}-2\varepsilon\langle x,a\rangle+\varepsilon^{2}b,\end{split}$
@@
Therefore, $0\leqslant D^{2}f_{\varepsilon}=D^{2}u_{\varepsilon}+\lambda I$ and the thesis follows. ∎
@@
$u$ is of class $\lambda$-$C^{1,1}$ (that is, $u\in C^{1,1}$ and the Lipschitz constant of $Du$ is $\lambda$) if and only if both ±$\pm u$ are $\lambda$-quasi-convex.
@@
Suppose $u$ is of class $\lambda$-$C^{1,1}$ and let $f\vcentcolon=u+\frac{\lambda}{2}|\cdot|^{2}$. We prove that $u$ is $\lambda$-quasi-convex; the proof for $-u$ is analogous. We have
@@
thus proving (cf. Proposition 1.13) that $f$ is convex, i.e. $u$ is $\lambda$-quasi-convex.
Conversely, suppose ±$\pm u$ are $\lambda$-quasi-convex. We first assume that $u$ is smooth; then we have $-\lambda I\leqslant D^{2}u(x)\leqslant\lambda I$ for all $x$. By the Mean Value Theorem,
@@
For the general case, note that quasi-convexity and quasi-concavity together imply $u$ being $C^{1}$. Indeed, both $f_{\pm}\vcentcolon=\pm u+\frac{\lambda}{2}|\cdot|^{2}$ have a supporting hyperplane from below at every point; this means that, for every $x$, there exist $p_{\pm}=p_{\pm}(x)$ such that the affine (and thus quadratic) functions $\varphi_{\pm}\vcentcolon=-f_{\pm}(x)-\langle p_{\pm},\cdot-x\rangle$ are upper contact functions for $-f_{\pm}$ at $x$ (in the sense of Lemma 1.21). Note now that $-f_{\pm}$ are both $2\lambda$-quasi-convex, since both ±$\pm u$ are $\lambda$-quasi-convex, and thus by Lemma 1.21 (cf. also Remark 2.3) we know that $p=D(-f_{+})(x)$, $q=D(-f_{-})(x)$, yielding $(q-p)/2=Du(x)$. That is to say, $u$ is differentiable everywhere. By Corollary 1.16 we conclude that $u$ is $C^{1}$.
Finally, approximate $u$ by $u_{\varepsilon}$ and recall that $Du_{\varepsilon}\to Du$ locally uniformly, since $u$ is $C^{1}$ (see, for instance, [49, Theorems 9.3 and 9.8]). Therefore, since by Lemma 2.4 and what we showed above in the smooth case we know that $Du_{\varepsilon}$ is $\lambda$-Lipschitz, so $Du$ is and we are done. ∎
@@ 2.3. Upper contact jets
The next definition is of crucial importance for all that follows. We recall that $\mathcal{S}(n)$ denotes the space of symmetric $n\times n$ real matrices which is equipped with its standard partial ordering (the Loewner order) where the quadratic form associated to $A\in\mathcal{S}(n)$ will often be normalized:
@@
Note that we always have equality in (2.4) for $y=x$ which justifies the use of contact point for $u$ at $x$. Also notice that $(p,A)\in J^{2,+}_{x}u$ if and only if the unique quadratic function $\varphi$ such that $J^{2}_{x}\varphi\vcentcolon=(\varphi,D\varphi,D^{2}\varphi)(x)=(u(x),p,A)$ satisfies (2.2); that is, if and only if $\varphi$ is a quadratic upper test function for $u$ at $x$ in the viscosity sense.
@@
$u(y)\leqslant u(x)+\langle p,y-x\rangle+Q_{A}(y-x)-\varepsilon|y-x|^{2}\qquad\forall y\in X\ \text{near}\ x;$ (2.5)
@@
Notice that the validity of inequality (2.3) for any upper test function $\varphi$ for $u$ at $x$, which defines the notion of $u$ being a subsolution of the equation (2.1): $F[u]=0$ at $x$, is equivalent to the validity, for $(p,A)=(D\varphi(x),D^{2}\varphi(x))$ for some such $\varphi$, of
@@
Notice (cf. Remark 2.7) that such jets are not upper contact jets in general, yet they are little-o upper contact jets; that is,
@@
Hence, we have five families of upper contact jets (for $u$ at $x$), namely
@@
    $J_{S}(x,u)$, $\varepsilon$-strict upper contact jets, for some $\varepsilon>0$ possibly different from jet to jet,
@@
    $J_{s}(x,u)$, strict upper contact jets according to Definition 2.6,
@@
    $J_{Q}(x,u)$, quadratic upper contact jets; that is, $J_{Q}(x,u)=J^{2,+}_{x}u$ is the set of all the upper contact jets for $u$ at $x$ according to Definition 2.6,
@@
    $J_{C^{2}}(x,u)$, $C^{2}$ upper contact jets associated to upper test functions $\varphi\in C^{2}$; the usual ones of viscosity theory,
@@
and one can prove (see [8, Lemma C.1]) that
@@
so that the closures of all sets in (2.8) in fact coincide. By the continuity of the operator $F$, this shows that, in order to prove that $u$ is a viscosity subsolution of (2.1) at $x$, one can check for (2.6) to hold for all upper contact jets in any of the families above.
@@
Suppose $u$ is twice differentiable at $x\in X$ (as open subset); then
@@
$u(y)=u(x)+\langle Du(x),y-x\rangle+Q_{D^{2}u(x)}(y-x)+o(|y-x|^{2})\quad\forall\,y\ \text{near}\ x.$ (2.9)
@@
Remark 2.10 (Lower contact functions and jets).
@@ 2.4. The theorems on upper contact jets and on summands
The main result on upper contact jets we wish to present can be found in [25, Theorem 4.1]. Two thirds of this result have been essentially proven above since these parts are the easy and natural extensions to quasi-convex functions of frist order properties already established for convex functions in LABEL:convediff. The third part concerning the partial upper semicontinuity of second derivatives will be “new” also per convex functions.
@@
Let $u$ be quasi-convex on a neighborhood of $x$ and suppose $(p,A)$ is an upper contact jet for $u$ at $x$. Then
  • (D at UCP)  if $(p,A)$ is an upper contact jet for $u$ at $x$, then $u$ is differentiable at $x$ and $p=Du(x)$ is unique.
Furthermore, for every set $E$ of full measure in a neighborhood of $x$ there exist a sequence $\{x_{j}\}_{j\in\mathbb{N}}\subset E\cap\operatorname{Diff}^{2}\!u$, with $x_{j}\to x$ as $j\to\infty$, and $\bar{A}\in\mathcal{S}(n)$ such that
  • (PC of FD)  $Du(x_{j})\to Du(x)=p$;
  • (PUSC of SD)  $D^{2}u(x_{j})\to\bar{A}\leqslant A$.
The proof of the theorem reduces to the proof of the three Lemmas 2.13, 2.14 and 2.15 below. As suggested above, the first two lemmas follow directly from corresponding results that we proved for convex functions. In particular, the differentiability at upper contact points of Lemma 2.13 is the extension to quasi-convex functions (reformulated in the language of upper contact jets) of the differentiability at points of upper quadratic contact for convex functions of Lemma 1.21 and the partial continuity of first derivatives of Lemma 2.14 for quasi-convex functions is immediately deduced from formula (1.18) of Corollary 1.16 which showed that convex differentiable functions are of class $C^{1}$. For the sake of completeness, we will include their short proofs.
@@
Let $u$ be locally quasi-convex. If $x$ is an upper contact point for $u$, then $u$ is differentiable at $x$. Moreover, if $(p,A)$ is any upper contact jet for $u$ at $x$, then $p=Du(x)$ is unique.
@@
There exists a ball $B$ about $x$ and a quadratic function $\varphi$ such that $u$ is $\lambda$-quasi-convex on $B$ for some $\lambda>0$ and $u\leqslant\varphi$ on $B$, with $u(x)=\varphi(x)$ (that is, $\varphi$ is an upper contact quadratic function for $u$ at $x$). Hence $\tilde{\varphi}\vcentcolon=\varphi+\frac{\lambda}{2}|\cdot|^{2}$ is an upper contact quadratic function for the convex function $\tilde{u}\vcentcolon=u+\frac{\lambda}{2}|\cdot|^{2}$ at $x$. Therefore by Lemma 1.21 $\tilde{u}$ is differentiable at $x$ with $D\tilde{u}(x)=D\tilde{\varphi}(x)$; that is, $u$ is differentiable at $x$ with $Du(x)=D\varphi(x)$. Finally, the same argument also proves that $p=Du(x)$ is unique for any $(p,A)\in J^{2,+}_{x}u$ (cf. also Proposition 2.9). ∎
@@
Let $u$ be locally quasi-convex, and $x_{j}\to x$. If $u$ is differentiable at each $x_{j}$ and at $x$, then $Du(x_{j})\to Du(x)$.
@@
Lemma 2.13 clearly gives the first point of Theorem 2.12, while Lemmas 2.13 and 2.14 together give the second point. Indeed, since $(p,A)\in J^{2,+}_{x}u$, by Lemma 2.13 $x\in\operatorname{Diff}^{1}\!u$, and by Rademacher’s and Alexandrov’s theorems $|{\operatorname{Diff}^{k}\!u\cap\mathcal{U}}|=\left|\,\mathcal{U}\right|$ for $k=1,2$, where we denoted by $\mathcal{U}$ the neighborhood of $x$ which we consider. Also, we know that $x$ is a limit point for $E\subset\mathcal{U}$ with $|E|=|\,\mathcal{U}|$ (otherwise there would exists a small ball $B_{\rho}(x)\subset\mathcal{U}\cap E^{\mathsf{c}}$, yielding $|E|<|\,\mathcal{U}|$, contradiction). Analogously $x$ is a limit point for $E\cap\operatorname{Diff}^{1}\!u$ as well, hence we can conclude by invoking Lemma 2.14.
@@
Let $u$ be a locally quasi-convex function, and $E$ be a set of full measure in a neighborhood of $x$. If $(p,A)$ is an upper contact jet for $u$ at $x$, then there exists a sequence of upper contact points $\{x_{j}\}\subset E\cap\operatorname{Diff}^{2}(u)$ such that $D^{2}u(x_{j})\to\bar{A}\in\mathcal{S}(n)$ with $\bar{A}\leqslant A$.
While the Upper Contact Jet Theorem 2.12 collects the basic information about upper contact jets of a single locally convex function $u$, the following result shows that for the sum of two locally quasi-convex functions one obtains useful information on the upper contact jets of the summands from each upper contact jet of the sum. Interestingly, this theorem, which is [26, Theorem 7.1], can be interpreted in multiple ways:
@@
    it obviously complements the partial upper semicontinuity of second derivatives for quasi-convex functions;
@@
    it easily yields a quasi-convex version of the so-called Theorem on Sums, which plays a crucial role in the viscosity theory;
@@
Suppose $u$ and $v$ are locally quasi-convex and that the sum $w\vcentcolon=u+v$ has an upper contact jet $(p,A)$ at $x$. Then the following hold:
@@
    $x$ is an upper contact point for both the summands $u$ and $v$ (which are therefore differentiable at $x$ by Lemma 2.13), whose upper contact jets at $x$ are of the form $(Du(x),\ast)$ and $(Dv(x),\ast)$;
@@
    for every set $E$ of full measure in a neighborhood of $x$, there exist a sequence $\{x_{j}\}_{j\in\mathbb{N}}$ of upper contact points for $w$, $x_{j}\in E\cap\operatorname{Diff}^{2}\!u\cap\operatorname{Diff}^{2}\!v$, $x_{j}\to x$, and $B,C\in\mathcal{S}(n)$, such that
[…]
@@
For the claim (ii), we begin by noting that the sum $w$ of locally quasi-convex functions $u$ and $v$ is clearly locally quasi-convex. Now, by applying Lemma 2.15 to the sum $w$ there exist a sequence $\{x_{j}\}\subset E\cap\operatorname{Diff}^{2}\!u\cap\operatorname{Diff}^{2}\!v$ and a matrix $\bar{A}\in\mathcal{S}(n)$, $\bar{A}\leqslant A$, such that $x_{j}\to x$ and $D^{2}w(x_{j})\to\bar{A}$ as $j\to\infty$. [^3] Note that in order to apply the lemma we are considering $E\cap\operatorname{Diff}^{2}(u)$ as the set of full measure near $x$. Finally, if
@@
then there exist $B,C\in\mathcal{S}(n)$ such that, up to a subsequence, $D^{2}u(x_{j})\to B$ and $D^{2}v(x_{j})\to C$, where $B+C=\lim_{j\to\infty}\left(D^{2}u(x_{j})+D^{2}v(x_{j})\right)=\lim_{j\to\infty}D^{2}w(x_{j})=\bar{A}\leqslant A$. In order to prove (2.11), let $\lambda>0$ be such that both $u$ and $v$ are $\lambda$-quasi-convex in a neighborhood of $x$, and note that for any $\varepsilon>0$ there exists $j_{0}=j_{0}(\varepsilon)\in\mathbb{N}$ such that
@@
We used the basic fact that each interval $[A,B]\subset(\mathcal{S}(n),\leqslant)$ is compact (with respect to the subspace topology inherited from ${\mathbb{R}}^{n^{2}}$). To prove it, let $Y\in[A,B]$ and note that by the Courant–Fischer min-max theorem one has that $\lambda_{i}(A)\leqslant\lambda_{i}(Y)\leqslant\lambda_{i}(B)$ for all $i=1,\dots,n$. Hence, if the eigenvalues are arranged in increasing order, the entries of $Y$ are bounded by $n\left(|\lambda_{1}(A)|\vee|\lambda_{n}(B)|\right)$, thus, since it is trivial that $[A,B]$ is closed, one concludes by invoking the Heine–Borel theorem.
As we said, a version of the Theorem on Sums for quasi-convex functions immediately follows if we double the variables, as it is easy to see.
Corollary 2.19 (Theorem on Sums for quasi-convex functions).
Let $u$ and $v$ be $\lambda$-quasi-convex on the open subsets $X$ and $Y$ of ${\mathbb{R}}^{n}$, respectively. Define
@@
By induction, one can also extend this result by adding additional variables, so that it assumes a form which more closely resembles the classical Theorem on Sums for semicontinuous functions, in the equivalent formulation one finds in [13, Appendix, Theorem ${}^{\prime}$]. We leave it to the reader to write an appropriate statement.

### v3

@@ 2.2. A semiconvexity characterization of $C^{1,1}$
$\eta_{\varepsilon}\vcentcolon=\frac{1}{\varepsilon^{n}}\,\eta\!\left(\frac{\cdot}{\varepsilon}\right),\quad\forall\varepsilon>0,$ (2.1)
so that the collection $\{\eta_{\varepsilon}\}_{\varepsilon>0}$ is be the so-called approximation of the identity based on $\eta$ (cf., e.g., [52, Section 9.2]).
@@
Suppose $u$ is $\lambda$-semiconvex on ${\mathbb{R}}^{n}$. Let $\eta$ be a mollifier, and let $\{\eta_{\varepsilon}\}_{\varepsilon>0}$ the approximation of the identity based on $\eta$. Then $u_{\varepsilon}\vcentcolon=u\ast\eta_{\varepsilon}$ is $\lambda$-semiconvex.
@@
Let $f\vcentcolon=u+\frac{\lambda}{2}|\cdot|^{2}$ be convex. It is well known that the mollification $f_{\varepsilon}\vcentcolon=f\ast\eta_{\varepsilon}$ is smooth and, moreover, is still convex. Indeed, since
$\begin{split}f_{\varepsilon}(tx+(1-t)z)&\vcentcolon=\int_{{\mathbb{R}}^{n}}f(tx+(1-t)z-y)\eta_{\varepsilon}(y)\,\mathrm{d}y\\ &=\int_{{\mathbb{R}}^{n}}f(t(x-y)+(1-t)(z-y))\eta_{\varepsilon}(y)\,\mathrm{d}y,\end{split}$
by the convexity of $f$ and the nonnegativity of $\eta$ one has
$f_{\varepsilon}(tx+(1-t)z)\leqslant\int_{{\mathbb{R}}^{n}}\bigl{(}tf(x-y)+(1-t)f(z-y)\bigr{)}\eta_{\varepsilon}(y)\,\mathrm{d}y\vcentcolon=tf_{\varepsilon}(x)+(1-t)f_{\varepsilon}(z).$
In addition, by making the change of variables $z=\varepsilon y$ and using that $\int_{{\mathbb{R}}^{n}}\eta=1$, one has
$\begin{split}\left(|\cdot|^{2}\ast\eta_{\varepsilon}\right)\!(x)\vcentcolon=\int_{{\mathbb{R}}^{n}}|x-z|^{2}\eta(z/\varepsilon)\varepsilon^{-n}\,\mathrm{d}z=|x|^{2}-2\varepsilon\langle x,a\rangle+\varepsilon^{2}b\vcentcolon=q_{\varepsilon}(x),\end{split}$
@@
Therefore, since $f_{\varepsilon}$ is smooth and convex
$0\leqslant D^{2}f_{\varepsilon}=D^{2}\Bigl{(}u_{\varepsilon}+\frac{\lambda}{2}q_{\varepsilon}\Bigr{)}=D^{2}u_{\varepsilon}+\lambda I,$
and the thesis follows from Proposition 1.28. ∎
@@
$u$ is of class $\lambda$-$C^{1,1}$ (that is, $u\in C^{1,1}$ and the Lipschitz constant of $Du$ is $\lambda$) if and only if both ±$\pm u$ are $\lambda$-semiconvex.
@@
Suppose $u$ is of class $\lambda$-$C^{1,1}$ and let $f\vcentcolon=u+\frac{\lambda}{2}|\cdot|^{2}$. We prove that $u$ is $\lambda$-semiconvex; the proof for $-u$ is analogous. We have
@@
thus proving (cf. Proposition 1.15) that $f$ is convex; that is, $u$ is $\lambda$-semiconvex.
Conversely, suppose ±$\pm u$ are $\lambda$-semiconvex. We first assume that $u$ is smooth; then we have $-\lambda I\leqslant D^{2}u(x)\leqslant\lambda I$ for all $x$. By the Mean Value Theorem,
@@
For the general case, note that semiconvexity and semiconcavity together imply $u$ being $C^{1}$. Indeed, both $f_{\pm}\vcentcolon=\pm u+\frac{\lambda}{2}|\cdot|^{2}$ have a supporting hyperplane from below at every point; this means that, for every $x$, there exist $p_{\pm}=p_{\pm}(x)$ such that the affine (and thus quadratic) functions $\varphi_{\pm}\vcentcolon=-f_{\pm}(x)-\langle p_{\pm},\cdot-x\rangle$ are upper contact functions for $-f_{\pm}$ at $x$ (in the sense of Lemma 1.23). Note now that $-f_{\pm}$ are both $2\lambda$-semiconvex, since both ±$\pm u$ are $\lambda$-semiconvex, and thus by Lemma 1.23 (cf. also Remark 2.3) we know that $p=D(-f_{+})(x)$, $q=D(-f_{-})(x)$, yielding $(q-p)/2=Du(x)$. That is to say, $u$ is differentiable everywhere. By Corollary 1.18 we conclude that $u$ is $C^{1}$.
Finally, approximate $u$ by $u_{\varepsilon}$ and recall that $Du_{\varepsilon}\to Du$ locally uniformly, since $u$ is $C^{1}$ (see, for instance, [52, Theorems 9.3 and 9.8]). Therefore, since by Lemma 2.4 and what we showed above in the smooth case we know that $Du_{\varepsilon}$ is $\lambda$-Lipschitz, so is $Du$ and we are done. ∎
@@ 2.3. Upper contact jets
The next definition is of crucial importance for all that follows. We recall that $\mathcal{S}(n)$ denotes the space of symmetric $n\times n$ real matrices which is equipped with its standard partial ordering (the Loewner order) where the quadratic form associated to $A\in\mathcal{S}(n)$ will often be normalized as
@@
Note that we always have equality in (2.2) for $y=x$ which justifies the use of contact point for $u$ at $x$. Also notice that $(p,A)\in J^{2,+}_{x}u$ if and only if the unique quadratic function $\varphi$ such that $J^{2}_{x}\varphi\vcentcolon=(\varphi,D\varphi,D^{2}\varphi)(x)=(u(x),p,A)$ satisfies
$u\leqslant\varphi\quad\text{near $x$, with equality at $x$,}$ (2.3)
that is, if and only if $\varphi$ is a quadratic upper test function for $u$ at $x$ in the viscosity sense.
@@
$u(y)\leqslant u(x)+\langle p,y-x\rangle+Q_{A}(y-x)-\varepsilon|y-x|^{2}\qquad\forall y\in X\ \text{near}\ x;$
@@
Notice that the validity of inequality $F[\varphi]\vcentcolon=F(x,\varphi(x),D\varphi(x),D^{2}\varphi(x))\geqslant 0$ for any upper test function $\varphi$ for $u$ at $x$, which defines the notion of $u$ being a subsolution of the equation $F[u]=0$ at $x$, is equivalent to the validity, for $(p,A)=(D\varphi(x),D^{2}\varphi(x))$ for some such $\varphi$, of
@@
Such jets are not upper contact jets in general, yet they are little-o upper contact jets; that is (cf. Remark 2.7),
@@
Hence, we have five families of upper contact jets (for $u$ at $x$):
@@
    $J_{S}(x,u)$, $\varepsilon$-strict upper contact jets, for some $\varepsilon>0$ possibly different from jet to jet;
@@
    $J_{s}(x,u)$, strict upper contact jets according to Definition 2.6;
@@
    $J_{Q}(x,u)$, quadratic upper contact jets; that is, $J_{Q}(x,u)=J^{2,+}_{x}u$ is the set of all the upper contact jets for $u$ at $x$ according to Definition 2.6;
@@
    $J_{C^{2}}(x,u)$, $C^{2}$ upper contact jets associated to upper test functions $\varphi\in C^{2}$; the usual ones of viscosity theory;
@@
and one can prove (see [9, Lemma C.1]) that
@@
so that the closures of all sets in (2.6) in fact coincide. By the continuity of the operator $F$, this shows that, in order to prove that $u$ is a viscosity subsolution of $F[u]=0$ at $x$, one can check for (2.4) to hold for all upper contact jets in any of the families above.
@@
Suppose $u$ is twice differentiable at $x\in X$ (open subset); then
@@
$u(y)=u(x)+\langle Du(x),y-x\rangle+Q_{D^{2}u(x)}(y-x)+o(|y-x|^{2})\quad\forall\,y\ \text{near}\ x.$
@@
Remark 2.10 (Lower contact points and jets).
@@ 2.4. The theorems on upper contact jets and on summands
The main result on upper contact jets we wish to present can be found in [28, Theorem 4.1]. Two thirds of this result have been essentially proven above since these parts are the easy and natural extensions to semiconvex functions of first order properties already established for convex functions in Chapter 1. The third part concerning the partial upper semicontinuity of second derivatives will be “new” also for convex functions.
@@
Let $u$ be semiconvex on a neighborhood of $x$ and suppose $(p,A)$ is an upper contact jet for $u$ at $x$. Then the following holds:
  • D at UCP, differentiability at upper contact points: if $(p,A)$ is an upper contact jet for $u$ at $x$, then $u$ is differentiable at $x$ and $p=Du(x)$ is unique.
Furthermore, for every set $E$ of full measure in a neighborhood of $x$ there exist a sequence $\{x_{j}\}_{j\in\mathbb{N}}\subset E\cap\operatorname{Diff}^{2}\!u$, with $x_{j}\to x$ as $j\to\infty$, and $\bar{A}\in\mathcal{S}(n)$ such that the following hold:
  • PC of FD, partial continuity of first derivatives: $Du(x_{j})\to Du(x)=p$;
  • PUSC of SD, partial upper semicontinuity of second derivatives: $D^{2}u(x_{j})\to\bar{A}\leqslant A$.
The proof of the theorem reduces to the proof of the three Lemmas 2.13, 2.14 and 2.15 below. As suggested above, the first two lemmas follow directly from corresponding results that we proved for convex functions. In particular, the differentiability at upper contact points of Lemma 2.13 is the extension to semiconvex functions (reformulated in the language of upper contact jets) of the differentiability at points of upper quadratic contact for convex functions of Lemma 1.23 and the partial continuity of first derivatives of Lemma 2.14 for semiconvex functions is immediately deduced from formula (1.14) of Corollary 1.18 which showed that convex differentiable functions are of class $C^{1}$. For the sake of completeness, we will include their short proofs.
@@
Let $u$ be locally semiconvex. If $x$ is an upper contact point for $u$, then $u$ is differentiable at $x$. Moreover, if $(p,A)$ is an upper contact jet for $u$ at $x$, then $p=Du(x)$ is unique.
@@
There exists a ball $B$ about $x$ and a quadratic function $\varphi$ such that $u$ is $\lambda$-semiconvex on $B$ for some $\lambda>0$ and $u\leqslant\varphi$ on $B$, with $u(x)=\varphi(x)$ (that is, $\varphi$ is an upper contact quadratic function for $u$ at $x$). Hence $\tilde{\varphi}\vcentcolon=\varphi+\frac{\lambda}{2}|\cdot|^{2}$ is an upper contact quadratic function for the convex function $\tilde{u}\vcentcolon=u+\frac{\lambda}{2}|\cdot|^{2}$ at $x$. Therefore by Lemma 1.23 $\tilde{u}$ is differentiable at $x$ with $D\tilde{u}(x)=D\tilde{\varphi}(x)$; that is, $u$ is differentiable at $x$ with $Du(x)=D\varphi(x)$. Finally, the same argument also proves that $p=Du(x)$ is unique for any $(p,A)\in J^{2,+}_{x}u$ (cf. also Proposition 2.9). ∎
@@
Let $u$ be locally semiconvex, and $x_{j}\to x$. If $u$ is differentiable at each $x_{j}$ and at $x$, then $Du(x_{j})\to Du(x)$.
@@
Lemma 2.13 clearly gives the first point of Theorem 2.12, while Lemmas 2.13 and 2.14 together give the second point. Indeed, since $(p,A)\in J^{2,+}_{x}u$, by Lemma 2.13, $x\in\operatorname{Diff}^{1}\!u$, and by Rademacher’s and Alexandrov’s theorems, denoting by $\mathcal{U}$ the neighborhood of $x$ which we consider, $|{\operatorname{Diff}^{k}\!u\cap\mathcal{U}}|=\left|\,\mathcal{U}\right|$ for $k=1,2$. Also, we know that $x$ is a limit point for $E\subset\mathcal{U}$ with $|E|=|\,\mathcal{U}|$ (otherwise there would exists a small ball $B_{\rho}(x)\subset\mathcal{U}\cap E^{\mathsf{c}}$, yielding $|E|<|\,\mathcal{U}|$, contradiction). Analogously $x$ is a limit point for $E\cap\operatorname{Diff}^{1}\!u$ as well, hence we can conclude by invoking Lemma 2.14.
@@
Let $u$ be a locally semiconvex function, and $E$ be a set of full measure in a neighborhood of $x$. If $(p,A)$ is an upper contact jet for $u$ at $x$, then there exists a sequence of upper contact points $\{x_{j}\}_{j\in\mathbb{N}}\subset E\cap\operatorname{Diff}^{2}(u)$ such that $D^{2}u(x_{j})\to\bar{A}\in\mathcal{S}(n)$ with $\bar{A}\leqslant A$.
While the Upper Contact Jet Theorem 2.12 collects the basic information about upper contact jets of a single locally convex function $u$, the following result shows that for the sum of two locally semiconvex functions one obtains useful information on the upper contact jets of the summands from each upper contact jet of the sum. Interestingly, this theorem, which is [29, Theorem 7.1], can be interpreted in multiple ways:
@@
    it obviously complements the partial upper semicontinuity of second derivatives for semiconvex functions;
@@
    it easily yields a semiconvex version of the so-called Theorem on Sums, which plays a crucial role in the viscosity theory;
@@
Suppose $u$ and $v$ are locally semiconvex and that the sum $w\vcentcolon=u+v$ has an upper contact jet $(p,A)$ at $x$. Then the following hold:
@@
    $x$ is an upper contact point for both the summands $u$ and $v$ (which are therefore differentiable at $x$ by Lemma 2.13), whose upper contact jets at $x$ are of the form $(Du(x),-)$ and $(Dv(x),-)$;
@@
    for every set $E$ of full measure in a neighborhood of $x$, there exist $B,C\in\mathcal{S}(n)$ and a sequence $\{x_{j}\}_{j\in\mathbb{N}}$ of upper contact points for $w$, with $x_{j}\in E\cap\operatorname{Diff}^{2}\!u\cap\operatorname{Diff}^{2}\!v$ and $x_{j}\to x$ as $j\to\infty$, such that
    $Du(x_{j})\to Du(x),\qquad Dv(x_{j})\to Dv(x),$
    $D^{2}u(x_{j})\to B,\quad D^{2}v(x_{j})\to C,\quad\text{with $B+C\leqslant A$}.$
@@
For the claim (ii), we begin by noting that the sum $w$ of locally semiconvex functions $u$ and $v$ is clearly locally semiconvex. Now, by applying Lemma 2.15 to the sum $w$ there exist a sequence $\{x_{j}\}\subset E\cap\operatorname{Diff}^{2}\!u\cap\operatorname{Diff}^{2}\!v$ and a matrix $\bar{A}\in\mathcal{S}(n)$, $\bar{A}\leqslant A$, such that $x_{j}\to x$ and $D^{2}w(x_{j})\to\bar{A}$ as $j\to\infty$. [^1] Note that to apply the lemma we are considering $E\cap\operatorname{Diff}^{2}(u)$ as the set of full measure near $x$. Finally, if
@@
then there exist $B,C\in\mathcal{S}(n)$ such that, up to a subsequence, $D^{2}u(x_{j})\to B$ and $D^{2}v(x_{j})\to C$, where $B+C=\lim_{j\to\infty}\left(D^{2}u(x_{j})+D^{2}v(x_{j})\right)=\lim_{j\to\infty}D^{2}w(x_{j})=\bar{A}\leqslant A$. In order to prove (2.8), let $\lambda>0$ be such that both $u$ and $v$ are $\lambda$-semiconvex in a neighborhood of $x$, and note that for any $\varepsilon>0$ there exists $j_{0}=j_{0}(\varepsilon)\in\mathbb{N}$ such that
@@
We used the basic fact that each interval $[A,B]\subset(\mathcal{S}(n),\leqslant)$ is compact (with respect to the subspace topology inherited from ${\mathbb{R}}^{n^{2}}$). To prove that, let $Y\in[A,B]$ and note that by the Courant–Fischer min-max theorem one has that $\lambda_{i}(A)\leqslant\lambda_{i}(Y)\leqslant\lambda_{i}(B)$ for all $i=1,\dots,n$. Hence, if the eigenvalues are arranged in increasing order, the entries of $Y$ are bounded by $n\left(|\lambda_{1}(A)|\vee|\lambda_{n}(B)|\right)$, thus, since it is trivial that $[A,B]$ is closed, one concludes by invoking the Heine–Borel theorem.
As we said, a version of the Theorem on Sums for semiconvex functions immediately follows if we double the variables, as it is easy to see.
Corollary 2.19 (Theorem on Sums for semiconvex functions).
Let $u$ and $v$ be $\lambda$-semiconvex on the open subsets $X$ and $Y$ of ${\mathbb{R}}^{n}$, respectively. Define
@@
By induction, one can also extend this result by adding additional variables, so that it assumes a form which more closely resembles the classical Theorem on Sums for semicontinuous functions, in the equivalent formulation one finds in [14, Appendix, Theorem 3.2]. We leave it to the reader to write an appropriate statement.